# Evo-0: Vision-Language-Action Model with Implicit Spatial Understanding

### v2

@@ I Introduction
However, existing VLA models exhibit a critical limitation: the lack of precise 3D spatial understanding. This limitation can be largely attributed to two main factors: (1) the pre-training data and objectives of VLMs are primarily based on 2D image-text alignments, and (2) the robotic datasets used for fine-tuning typically contain only RGB observations, lacking 3D spatial information. As a result, these models often struggle to capture the precise geometric and spatial relationships that are essential for effective interaction in the physical world. Recent studies [1, 2] have empirically validated this observation, showing that VLMs tend to generalize poorly when it comes to interpret 3D structures from visual inputs alone. This presents a critical bottleneck in scaling VLA models to more complex tasks that require accurate spatial reasoning and physical interaction.
To address this limitation, recent approaches [3, 4, 5, 6] have tried to incorporate 3D information into VLA models to enhance their spatial understanding capabilities. A common strategy is to explicitly inject depth information into the learning pipeline, such as point clouds or depth maps, either captured by depth sensors or predicted by pre-trained depth estimation networks [1, 7, 8]. While effective to some extent, these methods introduce new challenges. They often require additional depth sensors, which may not be available in practical settings. Moreover, the defective depth estimations can introduce extra noise, affecting the reliability of the learned 3D representations.
In this paper, we introduce Evo-0, a novel VLA architecture that explores an implicit strategy to enhance the spatial understanding of VLA models without relying on depth sensors or explicit depth estimation. We leverage a spatial encoder trained on large-scale 2D–3D paired data [9] to extract geometric features from RGB images, thereby enhancing the spatial understanding of VLA. To this end, we design a lightweight fusion module that integrates geometry features from spatial encoder with visual tokens in VLM, enabling the model to perceive object layouts and reason about spatial relations more effectively. We demonstrate the effectiveness of our method through comprehensive experiments on 5 simulation tasks, 5 real-world manipulation tasks, and a robustness evaluation with 5 distinct disturbance conditions. Across all settings, our model consistently enhances spatial understanding and outperforms state-of-the-art VLA models.
@@ II Related Work
Vision-Language-Action Models. Recently, several studies [10, 11, 12, 6, 13, 14] have focused on building general-purpose robot policies by extending pre-trained vision-language models (VLMs) with action prediction capabilities. These models, known as vision-language-action (VLA) models, demonstrate strong performance and few-shot generalization across a wide range of embodied tasks.
Among them, OpenVLA [10] is trained on 970k multi-robot demonstrations from the Open-X Embodiment [15] dataset, demonstrates strong generalization across a wide range of tasks and embodiments, and supports efficient fine-tuning under limited computational resources. $\pi_{0}$ [11] adapts the PaliGemma [16] architecture for robotic control and introduces a flow-matching-based [17, 18] action expert module that enables accurate prediction of continuous actions. GR00T [12] introduces an effective co-training strategy that jointly leverages web data, synthetic data, and real-world robot data within a unified framework, enabling broad generalization across tasks and embodiments.
@@
Robot Learning with 3D Information. In response to the spatial limitations of 2D-based VLA models, several recent approaches [1, 5, 3, 4, 19, 20, 21] have explored integrating 3D information to enhance spatial understanding. For example, 3D-VLA [5] fuses 3D perception, reasoning, and action through a 3D-based large language model [22], trained on a large-scale 3D dataset curated from existing embodied robotics benchmarks. To make 3D-aware policies applicable in real-world scenarios, methods such as SpatialVLA [4] and PointVLA [3] incorporate depth information captured from additional RGB-D cameras or depth estimation models, which enhances 3D scene understanding and enables more accurate perception of spatial relationships, object geometry, and depth-aware interactions.
@@
To address this issue, we propose integrating VGGT [9] into existing VLA models. While keeping the input as RGB images, VGGT implicitly models 3D structure by fusing spatial features from multi-view observations. Our approach serves as a bridge between pure 2D input models and explicit 3D perception methods, enhancing spatial understanding without requiring additional sensors or depth estimation modules.
@@ III-A Preliminaries
Vision-Language-Action Models. As a promising approach toward generalist robot policies, Vision-Language-Action (VLA) models have emerged as an increasingly popular research direction [11, 10, 3, 23]. VLAs aim to bridge the gap between high-level human instructions and low-level robotic actions by leveraging the rich multimodal priors encoded in large-scale pre-trained Vision-Language Models (VLMs), such as Paligemma [16], CLIP [24], LLaMA [25, 26], and Flamingo [27]. These VLMs are trained on vast and diverse internet-scale image-text pairs, endowing them with strong world knowledge and the ability to ground natural language in visual concepts.
@@
Visual Geometry Foundation Models. Unlike traditional SLAM or depth estimation pipelines that rely on finely-tuned modules and sensors, Visual Geometry Foundation Models (VGFMs) [28, 29, 9, 30] are a class of vision models trained to reconstruct 3D structural information from 2D visual inputs. Since VGFMs are trained with geometric supervision, they have the ability to recover fine-grained spatial structure from multi-view monocular inputs. These models provide strong structural priors for downstream tasks such as spatial understanding, especially when explicit 3D sensors are unavailable.
@@
Recently, Visual Geometry Grounded Transformer (VGGT) [9] has introduced a novel feed-forward architecture and demonstrated impressive performance in 3D attributes prediction. It takes an arbitrary number of image views as input and alternates between frame-wise and global self-attention to model spatial consistency. Given a sequence of $N$ RGB images $\{I^{i}\}_{i=1}^{N}$, where each $I_{i}\in\mathbb{R}^{3\times H\times W}$, the model outputs a set of 3D annotations for each frame, including predicted camera poses $g_{i}$, depth maps $D_{i}$, point maps $P_{i}$, and 3D point tracks $T_{i}$, i.e.,
@@ III-B Proposed VLA Architecture
Recent 3D-based VLA models, such as PointVLA [3] and SpatialVLA [4], often employ explicit 3D inputs like point clouds or depth maps to enhance spatial understanding. While effective, these approaches typically require additional sensors and preprocessing, and are often sensitive to variations in camera viewpoints. In contrast, VGGT presents a promising alternative for implicitly introducing spatial awareness, benefiting from its diverse training data and elegant feed-forward architecture. Recent studies have successfully applied VGGT to VLM architectures [31] and SLAM systems [32], demonstrating that geometry-grounded visual tokens can improve spatial understanding in both multimodal learning and classical robotic perception.
Motivated by these findings, we hypothesize that introducing geometry-aware visual representations from VGGT into the action prediction pipeline can enrich spatial context, leading to more precise and generalizable policy learning without requiring explicit point cloud or depth inputs. To evaluate this hypothesis, we build our model upon $\pi_{0}$ [11], a state-of-the-art open-source VLA model, and incorporate geometry-aware features from VGGT into its visual embedding stream. The architecture is described in Figure 1. Specifically, we utilize VGGT as a spatial encoder and extract tokens from its final layer:
@@
To integrate the VGGT-derived token features into the vision-language pipeline, we introduce a lightweight fuser module that combines embeddings from the Vision Transformer [33] and the VGGT encoder. Specifically, the fuser consists of a single cross-attention layer, where the 2D visual tokens $t_{2D}\in\mathbb{R}^{N\times M_{2D}\times d_{2D}}$ serve as queries, and the VGGT-derived tokens $t_{3D}\in\mathbb{R}^{N\times M_{3D}\times d_{3D}}$ act as keys and values. Here, $M_{2D}$ and $M_{3D}$ denote the number of tokens from ViT and VGGT encoder, respectively. The 2D visual tokens are then updated as follows:
@@
The fused tokens are then forwarded to the PaliGemma [16] vision-language model, which jointly attends over both the geometry-enhanced visual input and the language tokens to predict actions. To maintain computational efficiency and minimize disruption to the pretrained VLM backbone, we freeze the core VLM parameters and insert lightweight Low-Rank Adaptation (LoRA) [34] layers. During training, only the fuser module, LoRA layers, and the flow-matching action expert are fine-tuned, enabling effective adaptation with minimal overhead.
@@ IV-A Simulation Experiments
Implementation Details. Our framework builds upon the open-source VLA model [11]. We adopt OpenVLA-OFT [35] and $\pi_{0}$ as the baseline for comparison. The input comprises three RGB images, each resized to 224×224. Both the input state and the predicted output are represented as absolute joint angles. For a fair comparison, we utilize the official pretrained parameters provided by each method and adhere to their original training settings. For Evo-0, model is trained using the AdamW optimizer with a weight decay of $10^{-10}$ and a cosine learning rate schedule: the learning rate starts at $2.5\times 10^{-5}$, warms up over 1000 steps, and decays to $2.5\times 10^{-6}$. Training is conducted on a single NVIDIA A800 GPU (80 GB) with bfloat16 mixed precision, using a batch size of 32. For evaluation, we perform 20 rollouts for each task to assess success rates.

### v3

@@ I Introduction
However, existing VLA models exhibit a critical limitation: the lack of precise 3D spatial understanding. This limitation can be largely attributed to two main factors: (1) the pre-training data and objectives of VLMs are primarily based on 2D image-text alignments, and (2) the robotic datasets used for fine-tuning typically contain only RGB observations, lacking 3D spatial information. As a result, these models often struggle to capture the precise geometric and spatial relationships that are essential for effective interaction in the physical world. Recent studies [spatialbot, mm-spatial] have empirically validated this observation, showing that VLMs tend to generalize poorly when it comes to interpret 3D structures from visual inputs alone. This presents a critical bottleneck in scaling VLA models to more complex tasks that require accurate spatial reasoning and physical interaction.
To address this limitation, recent approaches [pointvla, spatialvla, 3d-vla, bridgevla] have tried to incorporate 3D information into VLA models to enhance their spatial understanding capabilities. A common strategy is to explicitly inject depth information into the learning pipeline, such as point clouds or depth maps, either captured by depth sensors or predicted by pre-trained depth estimation networks [spatialbot, bhat2023zoedepth, yang2024depth]. While effective to some extent, these methods introduce new challenges. They often require additional depth sensors, which may not be available in practical settings. Moreover, the defective depth estimations can introduce extra noise, affecting the reliability of the learned 3D representations.
In this paper, we introduce Evo-0, a novel VLA architecture that explores an implicit strategy to enhance the spatial understanding of VLA models without relying on depth sensors or explicit depth estimation. We leverage a spatial encoder trained on large-scale 2D–3D paired data [vggt] to extract geometric features from RGB images, thereby enhancing the spatial understanding of VLA. To this end, we design a lightweight fusion module that integrates geometry features from spatial encoder with visual tokens in VLM, enabling the model to perceive object layouts and reason about spatial relations more effectively. We demonstrate the effectiveness of our method through comprehensive experiments on 5 simulation tasks, 5 real-world manipulation tasks, and a robustness evaluation with 5 distinct disturbance conditions. Across all settings, our model consistently enhances spatial understanding and outperforms state-of-the-art VLA models.
@@ II Related Work
Vision-Language-Action Models. Recently, several studies [kim2024openvla, black2410pi0, bjorck2025gr00t, bridgevla, liu2024rdt, brohan2023rt] have focused on building general-purpose robot policies by extending pre-trained vision-language models (VLMs) with action prediction capabilities. These models, known as vision-language-action (VLA) models, demonstrate strong performance and few-shot generalization across a wide range of embodied tasks.
Among them, OpenVLA [kim2024openvla] is trained on 970k multi-robot demonstrations from the Open-X Embodiment [o2024open] dataset, demonstrates strong generalization across a wide range of tasks and embodiments, and supports efficient fine-tuning under limited computational resources. $\pi_{0}$ [black2410pi0] adapts the PaliGemma [beyer2024paligemma] architecture for robotic control and introduces a flow-matching-based [lipman2022flow, liu2022rectified] action expert module that enables accurate prediction of continuous actions. GR00T [bjorck2025gr00t] introduces an effective co-training strategy that jointly leverages web data, synthetic data, and real-world robot data within a unified framework, enabling broad generalization across tasks and embodiments.
@@
Robot Learning with 3D Information. In response to the spatial limitations of 2D-based VLA models, several recent approaches [spatialbot, 3d-vla, pointvla, spatialvla, chen2024sugar, goyal2024rvt, jia2024lift3d] have explored integrating 3D information to enhance spatial understanding. For example, 3D-VLA [3d-vla] fuses 3D perception, reasoning, and action through a 3D-based large language model [3d-llm], trained on a large-scale 3D dataset curated from existing embodied robotics benchmarks. To make 3D-aware policies applicable in real-world scenarios, methods such as SpatialVLA [spatialvla] and PointVLA [pointvla] incorporate depth information captured from additional RGB-D cameras or depth estimation models, which enhances 3D scene understanding and enables more accurate perception of spatial relationships, object geometry, and depth-aware interactions.
@@
To address this issue, we propose integrating VGGT [vggt] into existing VLA models. While keeping the input as RGB images, VGGT implicitly models 3D structure by fusing spatial features from multi-view observations. Our approach serves as a bridge between pure 2D input models and explicit 3D perception methods, enhancing spatial understanding without requiring additional sensors or depth estimation modules.
@@ III-A Preliminaries
Vision-Language-Action Models. As a promising approach toward generalist robot policies, Vision-Language-Action (VLA) models have emerged as an increasingly popular research direction [black2410pi0, kim2024openvla, pointvla, liu2025hybridvla]. VLAs aim to bridge the gap between high-level human instructions and low-level robotic actions by leveraging the rich multimodal priors encoded in large-scale pre-trained Vision-Language Models (VLMs), such as Paligemma [beyer2024paligemma], CLIP [radford2021learningtransferablevisualmodels], LLaMA [llama, llama2], and Flamingo [alayrac2022flamingovisuallanguagemodel]. These VLMs are trained on vast and diverse internet-scale image-text pairs, endowing them with strong world knowledge and the ability to ground natural language in visual concepts.
@@
Visual Geometry Foundation Models. Unlike traditional SLAM or depth estimation pipelines that rely on finely-tuned modules and sensors, Visual Geometry Foundation Models (VGFMs) [mast3r, Dust3r, vggt, megasam] are a class of vision models trained to reconstruct 3D structural information from 2D visual inputs. Since VGFMs are trained with geometric supervision, they have the ability to recover fine-grained spatial structure from multi-view monocular inputs. These models provide strong structural priors for downstream tasks such as spatial understanding, especially when explicit 3D sensors are unavailable.
@@
Recently, Visual Geometry Grounded Transformer (VGGT) [vggt] has introduced a novel feed-forward architecture and demonstrated impressive performance in 3D attributes prediction. It takes an arbitrary number of image views as input and alternates between frame-wise and global self-attention to model spatial consistency. Given a sequence of $N$ RGB images $\{I^{i}\}_{i=1}^{N}$, where each $I_{i}\in\mathbb{R}^{3\times H\times W}$, the model outputs a set of 3D annotations for each frame, including predicted camera poses $g_{i}$, depth maps $D_{i}$, point maps $P_{i}$, and 3D point tracks $T_{i}$, i.e.,
@@ III-B Proposed VLA Architecture
Recent 3D-based VLA models, such as PointVLA [pointvla] and SpatialVLA [spatialvla], often employ explicit 3D inputs like point clouds or depth maps to enhance spatial understanding. While effective, these approaches typically require additional sensors and preprocessing, and are often sensitive to variations in camera viewpoints. In contrast, VGGT presents a promising alternative for implicitly introducing spatial awareness, benefiting from its diverse training data and elegant feed-forward architecture. Recent studies have successfully applied VGGT to VLM architectures [wu2025spatialmllmboostingmllmcapabilities] and SLAM systems [maggio2025vggtslamdensergbslam], demonstrating that geometry-grounded visual tokens can improve spatial understanding in both multimodal learning and classical robotic perception.
Motivated by these findings, we hypothesize that introducing geometry-aware visual representations from VGGT into the action prediction pipeline can enrich spatial context, leading to more precise and generalizable policy learning without requiring explicit point cloud or depth inputs. To evaluate this hypothesis, we build our model upon $\pi_{0}$ [black2410pi0], a state-of-the-art open-source VLA model, and incorporate geometry-aware features from VGGT into its visual embedding stream. The architecture is described in Figure 1. Specifically, we utilize VGGT as a spatial encoder and extract tokens from its final layer:
@@
To integrate the VGGT-derived token features into the vision-language pipeline, we introduce a lightweight fuser module that combines embeddings from the Vision Transformer [dosovitskiy2020image] and the VGGT encoder. Specifically, the fuser consists of a single cross-attention layer, where the 2D visual tokens $t_{2D}\in\mathbb{R}^{N\times M_{2D}\times d_{2D}}$ serve as queries, and the VGGT-derived tokens $t_{3D}\in\mathbb{R}^{N\times M_{3D}\times d_{3D}}$ act as keys and values. Here, $M_{2D}$ and $M_{3D}$ denote the number of tokens from ViT and VGGT encoder, respectively. The 2D visual tokens are then updated as follows:
@@
The fused tokens are then forwarded to the PaliGemma [beyer2024paligemma] vision-language model, which jointly attends over both the geometry-enhanced visual input and the language tokens to predict actions. To maintain computational efficiency and minimize disruption to the pretrained VLM backbone, we freeze the core VLM parameters and insert lightweight Low-Rank Adaptation (LoRA) [lora] layers. During training, only the fuser module, LoRA layers, and the flow-matching action expert are fine-tuned, enabling effective adaptation with minimal overhead.
@@ IV-A Simulation Experiments
Implementation Details. Our framework builds upon the open-source VLA model [black2410pi0]. We adopt OpenVLA-OFT [kim2025fine] and $\pi_{0}$ as the baseline for comparison. The input comprises three RGB images, each resized to 224×224. Both the input state and the predicted output are represented as absolute joint angles. For a fair comparison, we utilize the official pretrained parameters provided by each method and adhere to their original training settings. For Evo-0, model is trained using the AdamW optimizer with a weight decay of $10^{-10}$ and a cosine learning rate schedule: the learning rate starts at $2.5\times 10^{-5}$, warms up over 1000 steps, and decays to $2.5\times 10^{-6}$. Training is conducted on a single NVIDIA A800 GPU (80 GB) with bfloat16 mixed precision, using a batch size of 32. For evaluation, we perform 20 rollouts for each task to assess success rates.